OpenStudy (anonymous):

How to do this : sqrt(A) + sqrt(B) + 2sqrt(C-3) + sqrt(D) + sqrt(E) = A+B+C+D+E+1. Determine the value of A*B*C*D*E

OpenStudy (anonymous):

Are there answer options?

OpenStudy (anonymous):

What? Don't you need 5 equations to solve 5 variables?

OpenStudy (anonymous):

no options, and no information what's kind number for A,B,C,D, and E. but i guess they are integer numbers.

OpenStudy (anonymous):

@NewOne luckily 5 variables aren't being solved for...

OpenStudy (abb0t):

You can combine them all to form one single root.

OpenStudy (anonymous):

@tanjung they can't be integers, sorry.

OpenStudy (abb0t):

Oh wait this is addition, sorry. No you can't.

OpenStudy (anonymous):

@tanjung are you sure it's addition?

OpenStudy (anonymous):

yes.. i think there is the special way to solve this, but i dont know.. it is an olympiad problem :)

OpenStudy (abb0t):

I think you can just set them each equal to each other and solve for each variable. I didn't read the full question. I thought it was = 0

OpenStudy (anonymous):

@abb0t if it were equal to 0 and you were interested in only real solutions, then A*B*C*D*E = 0 as well :-)

OpenStudy (abb0t):

Yeah, but they are = A + B + C ... 1

OpenStudy (anonymous):

any idea guys ? @AccessDenied , @experimentX , @Callisto ,@ganeshie8 @ghazi

OpenStudy (experimentx):

just put up 111 for 4 for c

OpenStudy (queelius):

Since we have no constraints, make a, b, d, and e equal 0. Then, we solve 2sqrt(c-3) = c + 1

OpenStudy (queelius):

This gives an imaginary solution. If you need the solution to be a real number, then you'll have to rethinking which values you choose for a, b, ...

OpenStudy (raden):

i think the last term should be -1, not +1... so that the equation can be a perfect square and to get a real solution!

OpenStudy (raden):

i assumed the last term is -1, so A+B+C+D+E-1=sqrt(A)+sqrt(B)+2sqrt(C-3)+sqrt(D)+sqrt(E) A-sqrt(A)+B-sqrt(B)+C-sqrt(C-3)+D-sqrt(D)+E-sqrt(E)-1=0 (sqrt(A)-1/2)^2 - 1/4 + (sqrt(B)-1/2)^2 - 1/4 + (sqrt(C-3)-1)^2 + 2 + (sqrt(D)-1/2)^2 - 1/4 + (sqrt(E)-1/2)^2 - 1/4 - 1 = 0 (sqrt(A)-1/2)^2 + (sqrt(B)-1/2)^2 + (sqrt(C-3)-1)^2 + (sqrt(D)-1/2)^2 + (sqrt(E)-1/2)^2 = 0 it must satisfies if all perfect squares in LHS will be zero. therefore, sqrt(A) = 1/2 -----> A=1/4 sqrt(B) = 1/2 -----> B=1/4 sqrt(C-3)=1 ------->C=4 sqrt(D) = 1/2 -----> D=1/4 sqrt(D) = 1/2 -----> E=1/4 so, the value of A*B*C*D*E = ....

OpenStudy (queelius):

Good catch, RadEn. I don't see how it's possible to get a real solution otherwise. I was just trying to solve 4*sqrt(x) = 3 + 4x... If instead of +1 we have -1, then it's 4sqrt(x) = 2+4x, which does have a real solution. (a=b=d=e=1/4)

OpenStudy (raden):

opsss.. the last shoulde be sqrt(E) = 1/2 -----> E=1/4, :)

OpenStudy (anonymous):

sqrt(A) <A , sqrt(B) <B, 2sqrt(C-3) <C , sqrt(D)<D , sqrt(E)<E so, sqrt(A) + sqrt(B) + 2sqrt(C-3) + sqrt(D) + sqrt(E)<A + B + C + D + E

OpenStudy (anonymous):

so, I guess it must be -1 as @RadEn said

OpenStudy (queelius):

Btw, saurav, sqrt(A) is greater than A if A < 1

OpenStudy (anonymous):

@sauravshakya your premises are faulty if they're in (0,1)

OpenStudy (anonymous):

But all variables cannot be 0 or 1

OpenStudy (anonymous):

So, if only one variale is non zero and it is not 1 then it must be sqrt(A) + sqrt(B) + 2sqrt(C-3) + sqrt(D) + sqrt(E)<A + B + C + D + E

OpenStudy (anonymous):

@sauravshakya that's not what (0,1) means... http://en.wikipedia.org/wiki/Interval_(mathematics)#Notations_for_intervals

OpenStudy (queelius):

saur, to make this problem have a real solution, start off by making c = 3. So now we have sqrt(A) + ... + sqrt(E) = A + B + D + E + 4. Assume A=B=D=E, then we have 4*sqrt(x) = 4x + 4. This has no real solution, but if we take Rad's advice, and make the +1 a -1, then it does.

OpenStudy (anonymous):

I was also trying to say the same thing

OpenStudy (anonymous):

to have a real solution it must be -1

OpenStudy (queelius):

*nods*